An extension cord made of two wires of diameter 0.129 cm (no. 16 copper wire with...

  1. An extension cord made of two wires of diameter 0.129 cm (no. 16 copper wire with r = 1.68x10-8) and of length 2.7 m (9 ft) is connected to an electric heater which draws 15.0 A on a 120-V line. How much power is dissipated in the cord? Please no hand written answers
Answer #1

P=IV,

P = IV

R = ρL / A

P = V^2/R

. First I found the power required by the heater. P = (15.0 A)(120 V) P = 1800 W

I wanted to find the power within the cords. R = ρL / A

A==4.16*10^-7*3.14

R=1.68*10^-8(2.7 m) / (4.16 * 10^-7)^2*3.14=



P = V^2 / R=

.